Prize Mustard uses a process costing system with two departments: (a) a Mixing Department, and (b) a Bottling Department. During May, Prize generated $120,000 in revenue by producing 40,000 gallons (10,000 cases) of mustard. The companyreduces its work in process inventories to zero eachmonth. Manufacturing costs for May are as follows Mixing Bottling Department Department Direct materials Direct labor Manufacturing overhea $16,000 $5,600 $11,400 $6,000 S3,900 $8,100 1 Refer to the above data. The unit cost per gallon of mustard processed by the Mixing Department in May was $ 2 Refer to the above data. The unit cost per case of mustard incurred by the Bottling Department in May was $ Refer to the above data. The total cost ofthe 10,000 cases transferred to finished goods in May was S 3 4 Refer to the above data. Assume that 9,500 cases of mustard were sold in May The amount charged to cost of goods sold in May S 5 Refer to the above data. Assume that $9,000 of finished goods were on hand at May 1. Under the LIFO method of inventory valuation, finished goods inventory at May 31 was S if 9.500 cases of mustard were sold in May
Answer #1

1. Total cost of Mixing department = $16,000 + $5,600 + $ 11,400 = $33,000

Unit cost per gallon of Mixing department is $33,000 / 40,000 = $ 0.825 per gallon

2.Total cost of bottling department = $6,000 + $3,900 + $8,100 = $18,000

Unit cost per case of mustard of Mixing department is $18,000 / 10,000 = $ 1.80 per case

3. Total cost of 10,000 cases of mustard is $33,000 + $18,000 = $51,000

4. Assume 9,500 cases of mustard were sold. The amount charged to cost of goods sold will be $51,000 / 10,000 * 9,500 = $48,450

5. Stock in Hand on May 1 is of $9,000,

Sale price per case = $120,000 / 10,000 = $120

So, the opening units were $9,000 / 120 = 75

Sold during May are 9,500 cases and produced are 10,000 cases. So the number of cases in the closing stock are 75 + 10,000 – 9,500 = 575

So, the finished goods inventory is $51,000 / 10,000 * 575 = $2,932.50

It is assumed that the opening finished goods are valued at sales price.